in a club, the ratio of boys to girls is 7 : 8. The ratio of one less than the number of girls to the total number of members is 1 : 2. How many boys and how many girls are in the club?
Answer by edjones(8007)

Let b=number of boys and g=number of girls.
b/g=7/8
8b=7g
.
(g-1)/b+g)=1/2
2g-2=b+g
b=g-2
.
(g-2)/g=7/8
8g-16=7g
g=16
b=14
.
Ed